# MGE-LDM: Joint Latent Diffusion for Simultaneous Music Generation and Source Extraction

### v1

@@ E.2 Additional Design Ablations
Table 7: Source extraction performance under different architectural and CFG settings. Metrics are reported as Log-Mel L1 distance $\downarrow$. All models are trained on $\textbf{S}_{A}\,$++\,+$\textbf{S}_{B}\,$. GN and LN denote GroupNorm and LayerNorm, respectively. $p$ indicates the classifier-free guidance (CFG) dropout rate applied to each track’s conditioning vector, and $s$ refers to the CFG guidance scale.
@@
We further examine how modeling and training design choices–such as normalization layers, classifier-free guidance (CFG) dropout probability, and CFG scale–affect extraction performance, and report the results in Table 7. When comparing normalization strategies, GroupNorm consistently matches or outperforms LayerNorm across most stems, demonstrating the advantage of modeling track-wise statistics in our multi-track architecture. This observation aligns with trends seen in mixture generation results. For CFG dropout, a higher dropout probability ($p=0.5$) leads to improved performance compared to the default $p=0.1$, suggesting that stronger stochastic conditioning is beneficial during source extraction. While this differs from the trend observed in mixture generation (Table 6), the discrepancy may be explained by the fact that, in extraction, non-target tracks are effectively treated as unconditioned. This makes overall performance more sensitive to the model’s ability to generalize in the presence of dropout. We also evaluate various CFG scales (1, 2, 4, 8). A scale of 1 yields the best performance overall, although scales 2 and 4 remain competitive. Performance degrades at scale 8, indicating that overly strong guidance can impair extraction quality.
@@ E.3 Iterative Generation Variants
Table 8: Preliminary results for iterative stem-wise generation. Metrics are reported as FAD $\downarrow$. Evaluation is conducted using a model trained exclusively on $\textbf{S}_{A}\,$. Each column header indicates the generation order of stems (e.g., BDGP denotes $\rightarrow$ $\rightarrow$ $\rightarrow$ piano).
@@
Appendix F Limitations
While MGE-LDM provides a flexible, class-agnostic framework for multi-track audio modeling, several limitations remain. First, all experiments are conducted using 16 kHz monaural audio, which restricts upper-frequency resolution and omits spatial information, limiting realism for high-fidelity or stereo music applications. Second, the model relies on CLAP-based semantic conditioning, which introduces a modality gap between text and audio. This can occasionally result in semantic drift during extraction, such as generating irrelevant instruments or hallucinating sources not present in the mixture, particularly for stems with limited training data. Third, although MGE-LDM reduces dependence on fixed instrument classes, it still requires multi-stem supervision during training. This restricts applicability to entirely unlabeled or web-scale audio collections.
Lastly, we observed that training on the MUSDB18 dataset alone with the same number of iterations as other configurations led to overfitting, likely due to the limited duration ($\sim$10 hours) of the training split. This highlights a challenge in low-resource multi-track settings.
@@ Appendix G Future Work
Future extensions of MGE-LDM include scaling to higher-resolution formats such as 44.1 kHz stereo audio, enabling richer timbral detail and spatial fidelity. To reduce the modality gap in text-conditioned extraction, fine-tuning on curated audio–text datasets like MusicCaps [7] is a promising direction. Given its minimal reliance on precise stem boundaries, MGE-LDM is naturally suited for incorporating weakly or noisily labeled multi-track data [95, 38], which may expand training diversity.
Another promising avenue is to pre-train MGE-LDM on large-scale mixture-only corpora–such as MTG-Jamendo [96] or the Free Music Archive [97]–to learn general audio priors for mixture track, followed by fine-tuning on multi-track datasets for source-aware generation. This two-stage training strategy is expected to enhance generative quality and improve generalization.
@@ Appendix H Spectrogram Examples of Generated Samples
We present Mel-spectrogram visualizations of generated audio samples across the three primary tasks: total generation, partial generation (imputation), and source extraction. All examples are produced by MGE-LDM trained on the combined Slakh2100 ($\textbf{S}_{A}\,$$+$$\textbf{S}_{B}\,$), MUSDB18 ($\textbf{M}_{u}\,$), and MoisesDB ($\textbf{M}_{o}\,$) datasets.

### v2

@@ E.2 Additional Design Ablations
Table 7: Source extraction performance under different architectural and CFG settings. Metrics are reported as Log-Mel L1 distance $\downarrow$. All models are trained on $\textbf{S}_{A}$$\textbf{S}_{B}$. GN and LN denote GroupNorm and LayerNorm, respectively. $p$ indicates the classifier-free guidance (CFG) dropout rate applied to each track’s conditioning vector, and $s$ refers to the CFG guidance scale.
@@
We further examine how modeling and training design choices, such as normalization layers, classifier-free guidance (CFG) dropout probability, and CFG scale, affect extraction performance and report the results in Table 7. When comparing normalization strategies, GroupNorm consistently matches or outperforms LayerNorm across most stems, demonstrating the advantage of modeling track-wise statistics in our multi-track architecture. This observation aligns with the trends seen in mixture generation results. For CFG dropout, a higher dropout probability ($p=0.5$) leads to improved performance compared to the default $p=0.1$, suggesting that stronger stochastic conditioning is beneficial during source extraction. While this differs from the trend observed in mixture generation (Table 6), the discrepancy may be explained by the fact that, in extraction, non-target tracks are effectively treated as unconditioned. This makes overall performance more sensitive to the model’s ability to generalize in the presence of dropout. We also evaluate various CFG scales (1, 2, 4, 8). A scale of 1 yields the best performance overall, although scales 2 and 4 remain competitive. Performance degrades at scale 8, indicating that overly strong guidance can impair extraction quality.
@@ E.3 Iterative Generation Variants
[…]
@@
Appendix F Extended Evaluation with Alternative Metrics
Table 9: Partial generation results evaluated with COCOLA [105] score $\uparrow$. Higher values indicate stronger coherence between the given mixture and the generated stems, complementing sub-FAD by emphasizing musical consistency. "Ground truth" values corresponmd to the COCOLA scores computed between the given mixture and the ground-truth accompaniment from the dataset. Bold values indicate the best results in each column (excluding Ground truth), and underlined values denote the best results among models trained on the same $\textbf{S}_{A}$ set.
Model Train Set $\textbf{S}_{A}$ $\textbf{S}_{B}$
MSDM $\times$ $\times$ $\times$ 57.74 51.87 57.76 56.10 45.96 57.64 54.48 50.67 49.40 55.75 43.25 40.46 52.82 47.09 - - - - - - - -
MSG-LD $\times$ $\times$ $\times$ 58.18 50.58 58.55 58.30 46.12 57.63 56.29 49.11 47.51 57.94 43.70 41.33 54.19 44.98 - - - - - - - -
MGE (ours) $\mathcal{T}_{1}$ $\times$ $\times$ $\times$ 60.93 53.71 58.09 59.78 49.77 60.43 59.16 52.95 51.90 60.03 46.28 43.46 56.77 49.73 63.20 56.44 59.52 60.59 62.35 61.11 63.74 58.66
$\mathcal{T}_{2}$ $\times$ $\times$ 61.29 51.43 62.61 61.35 48.79 60.94 59.12 50.82 49.39 60.21 46.06 42.81 56.73 47.01 64.32 65.21 65.46 62.31 64.31 62.33 64.52 61.10
$\mathcal{T}_{3}$ $\times$ $\times$ 55.40 44.76 57.29 59.34 43.90 55.18 56.09 45.68 45.42 55.91 44.10 43.35 51.44 43.53 60.23 54.74 56.70 60.18 56.50 60.66 62.03 61.00
$\mathcal{T}_{4}$ 56.73 49.65 60.58 61.06 46.92 57.88 57.82 50.66 49.93 58.32 46.53 45.39 54.02 46.86 65.94 64.10 64.47 63.77 59.49 62.39 65.02 58.52
Ground truth 60.11 53.55 56.33 57.16 50.42 59.74 58.59 53.11 52.29 57.80 47.06 44.32 56.07 50.18 64.58 64.34 64.85 63.38 64.75 63.20 65.29 62.16
To complement the quantitative results presented in Section 5, we provide additional evaluations using alternative metrics that better capture musical coherence and perceptual quality. While sub-FAD is widely adopted in accompaniment generation research [63, 34, 35, 36, 76, 81], it is considered suboptimal for evaluating musical coherence, since it primarily measures global timbral similarity rather than the harmonic and rhythmic interplay between stems [105]. To address this limitation, Ciranni et al. [105] introduced the COCOLA metric, which quantifies harmonic and rhythmic coherence through contrastive learning of musical audio representations.
We report the COCOLA results in Table 9 to complement our main findings. The metric is computed between the given mixture and the generated stems, and we additionally report the “Ground truth” values computed between the same mixture and the corresponding reference accompaniment from the dataset. Among models trained on $\textbf{S}_{A}$, our $\mathcal{T}_{1}$ variant outperforms the baselines on most instrument categories, except for guitar generation. Larger-scale models ($\mathcal{T}_{2}$$\mathcal{T}_{4}$) show similarly strong or even superior coherence, in several cases approaching or surpassing ground-truth levels. Since COCOLA emphasizes musical coherence rather than timbral accuracy with respect to reference stems, it provides a meaningful complement to sub-FAD, offering a more balanced view of partial generation quality.
[…]
@@ Appendix G Future Work
Future extensions of MGE-LDM include scaling to higher-resolution formats such as 44.1 kHz stereo audio, enabling richer timbral detail and spatial fidelity. In particular, this can be achieved by leveraging high-quality latent representations recently developed for the music domain [109, 110]. To reduce the modality gap in text-conditioned extraction, fine-tuning on curated audio–text datasets like MusicCaps [7] is a promising direction. Given its minimal reliance on precise stem boundaries, MGE-LDM is naturally suited for incorporating weakly or noisily labeled multi-track data [111, 39], which may expand training diversity.
Another promising avenue is to pre-train MGE-LDM on large-scale mixture-only corpora such as MTG-Jamendo [112] or the Free Music Archive [113] to learn general audio priors for mixture tracks, followed by fine-tuning on multi-track datasets for source-aware generation. This two-stage training strategy is expected to enhance generative quality and improve generalization.
@@
Table 10: Total generation results evaluated with $\text{FAD}_{\textit{CLAP-MA}}$ $\downarrow$.
Model Train Set Test Set
MSDM $\times$ $\times$ $\times$ .329 .290 .767 .744
MSG-LD $\times$ $\times$ $\times$ .191 .312 .658 .697
MGE (ours) $\mathcal{T}_{1}$ $\times$ $\times$ $\times$ .110 (.450) .285 .637 .627
$\mathcal{T}_{2}$ $\times$ $\times$ .368 (.359) .099 .578 .621
$\mathcal{T}_{3}$ $\times$ $\times$ .731 (.669) .632 .222 .165
$\mathcal{T}_{4}$ .563 (.712) .426 .235 .186
[…]
Model Train Set $\textbf{S}_{A}$ $\textbf{S}_{B}$
MSDM $\times$ $\times$ $\times$ .103 .100 .066 .096 .227 .189 .252 .135 .191 .302 .275 .357 .555 .403 - - - - - - - -
MSG-LD $\times$ $\times$ $\times$ .071 .070 .076 .077 .130 .135 .143 .117 .121 .157 .185 .188 .241 .190 - - - - - - - -
MGE (ours) $\mathcal{T}_{1}$ $\times$ $\times$ $\times$ .140 .170 .153 .154 .184 .175 .181 .200 .189 .184 .221 .218 .190 .214 .278 .086 .069 .327 .462 .201 .335 .118
$\mathcal{T}_{2}$ $\times$ $\times$ .259 .329 .251 .264 .420 .302 .362 .380 .361 .337 .500 .517 .422 .460 .425 .071 .129 .327 .330 .296 .123 .313
$\mathcal{T}_{3}$ $\times$ $\times$ .178 .202 .356 .253 .307 .402 .340 .368 .320 .477 .445 .469 .632 .484 .361 .087 .062 .045 .090 .087 .079 .060
$\mathcal{T}_{4}$ .170 .187 .367 .251 .253 .405 .308 .350 .289 .455 .450 .436 .571 .486 .312 .072 .107 .057 .340 .208 .085 .063
@@ Appendix H Spectrogram Examples of Generated Samples
We present Mel-spectrogram visualizations of generated audio samples across the three primary tasks: total generation, partial generation (imputation), and source extraction. All examples are produced by MGE-LDM trained on the combined Slakh2100 ($\textbf{S}_{A}$$\textbf{S}_{B}$), MUSDB18 ($\textbf{M}_{u}$), and MoisesDB ($\textbf{M}_{o}$) datasets.